Please help me to solve this equation 7n-2g=1,g=3n
Answer by nerdybill(7384)

The problem gives you two equations:
7n-2g=1 (equation 1)
g=3n (equation 2)
.
Starting with equation 1:
7n-2g=1
Substitute the definition of g from equation 2:
7n-2(3n)=1
7n-6n=1
n = 1
.
To find g, start with equation 2:
g = 3n = 3(1) = 3
.
Solution:
g = 3
n = 1